# VisFly: An Efficient and Versatile Simulator for Training Vision-based Flight

### v3

@@ IV Training examples
To validate this simulator, we present three cases of training vision-based flight policies: (1) learning to land, (2) learning to navigate in a cluttered environment, and (3) learning to cross the narrow gap cooperatively. We use reinforcement learning to train end-to-end models for controlling the drones using the vision inputs. To focus on training efficiency rather than the algorithms themselves, we employ PPO [38] from stable-baselines3[39] (SB3) for its outstanding capability to handle high-dimensional observationsThe backbone architecture of the policy network follows the standard design in SB3 , as shown in Figure 4, with modifications to enhance its adaptability across various tasks.
The image feature extractor is a conventional neural network that comprises three layers. Each layer includes a convolutional layer, an activation function, and a max pooling operation, followed by a linear layer with 128 units. The convolutional layers have kernel sizes of 5, 3, and 3, with channel counts of 6, 12, and 18. Each max pooling operation uses a kernel size of 2. The target extractor, state extractor, and swarm extractor are all multilayer perceptrons (MLPs) with two linear layers, having 128 and 64 units. The action net and value net are also MLPs, with two linear layers of 64 units. All activation functions are ReLU. This architecture has been parameterized, making it convenient to fine-tune. The action type in all the examples is CTBR. The quadrotor takes its full state (position, orientation, linear velocity, angular velocity), visual image, and target as observations. Most of the training hyper-parameters are set to default values. Detailed settings, the reward, done, and observation functions are provided in the source code.
Figure 4: Neural network architecture used across different tasks. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[40], MobileNet[41], EfficientNet[42]. The policies for the three tasks follow this architecture but are slightly different. VisFly additionally incorporates a recurrent network interface [43] for training tasks that do not strictly conform to the Markov process.
IV-A Learning to land
In this task, the quadrotor starts in flight, gradually decelerates, and eventually lands safely on the landing area. It continuously monitors the limited landing area directly below, which is represented as a black rectangle positioned horizontally on the ground within a clear garage, as shown in Figure 5. The quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in its field of view, strives to keep it centered, and incorporates it into its observation. During the flight, the quadrotor captures only $64\times 64$ RGB image from the camera rather than acquiring precise spatial information about the landing area. The center of the landing area is determined using threshold segmentation and included as a target observation. This training process takes 1.7 hours, with additional time required for prepossessing the center of the landing area, as detailed in Section IV-B.
Figure 5: Training results of landing, navigation, and swarm crossing. These subfigure columns respectively represent the results of landing, navigation, swarm crossing, including rendered image, reward curve and success rate curve. The final success rates of all tasks are approaching 100% approximately.
IV-B Learning to navigate in a cluttered environment
@@
IV-C Learning to cross a narrow gap cooperatively
We inserted two walls in the clear garage to create a narrow gap between them as Figure 5. Three quadrotors start from one side of the walls, fly towards distributed targets on the other side, and pass through the narrow gap simultaneously. They must cooperate by considering each other’s state to avoid obstacles and other agents safely. Each quadrotor uses a $64\times 64$ depth image as its visual observation, the final position as its target observation, and the states of other agents as its swarm observation. Unlike single-agent tasks, VisFly supports only one swarm per scene, which limits the maximum batch size due to memory constraints. Training takes 1.1 hours. The results show that by observing the states of other agents and the environment, the quadrotors can effectively fly across the gap and avoid collisions with each other.
V Discussion
VisFly’s PyTorch-based dynamics provide differentiable physics, though with slightly lower accuracy compared to professional simulators like Gazebo. However, lower accuracy does not impair the Sim-to-Real deployment. This has been demonstrated in challenging scenarios such as quadrotor racing [20] where the quadrotor is capable of flying close to its dynamic limits. Additionally, VisFly does not incorporate high-level controllers like Model Predictive Control (MPC) for two main reasons: First, these advanced controllers require substantial computational resources during simulation and real flight. Second, End-to-end learning is preferred because it closely mimics natural decision-making processes. In future work, we plan to integrate real-world dynamics into VisFly to enable Hardware-in-the-Loop (HITL) capabilities.
VI Conclusion

### v4

@@ III-D Domain randomization for sensor data and flight states
Figure 5: Neural network architecture used across different tasks. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[36], MobileNet[37], EfficientNet[38]. The policies for the three tasks follow this architecture but are slightly different. VisFly additionally incorporates a recurrent network interface [39]. Detailed policy setting is introduced in VisFly homepage.
Figure 6: Visualization of Left: navigation, Middle: cooperatively crossing, Right: landing tasks.
@@ IV Training examples
To validate this simulator, we present three cases of training vision-based flight policies: (1) learning to land, (2) learning to navigate in a cluttered environment, and (3) learning to cross the narrow gap cooperatively.
We use reinforcement learning to train end-to-end models for controlling the drones using vision inputs. To focus on training efficiency rather than the algorithms themselves, we employ PPO [40] from stable-baselines3[41] (SB3) for its outstanding capability to handle high-dimensional observations. The backbone architecture of the policy network follows the standard design in SB3, as shown in Figure 5, with modifications to enhance its adaptability across various tasks. The action type in all the examples is CTBR. The quadrotor takes its full state (position, orientation, linear velocity, angular velocity), visual image, and target as observations. Most of the hyper-parameters of PPO are set to default values. Detailed settings, the network architecture, the reward, done, and observation functions are provided in the source code.
Figure 7: Reward and success rate curves during training of navigation, corporately crossing, landing task.
IV-A Learning to navigate in a cluttered environment
@@
IV-B Learning to cross a narrow gap cooperatively
We inserted two walls in the clear garage to create a narrow gap between them as Figure 6. Three quadrotors start from one side of the walls, fly towards distributed targets on the other side, and pass through the narrow gap simultaneously. They must cooperate by considering each other’s state to avoid obstacles and other agents safely. Each quadrotor uses a $64\times 64$ depth image as its visual observation, the final position as its target observation, and the states of other agents as its group member observation. Unlike single-agent tasks, VisFly supports only one swarm per scene, which limits the maximum batch size due to memory constraints. This training takes 1.1 hours. The results show that by observing the states of other agents and the environment, the quadrotors can effectively fly across the gap and avoid collisions with each other.
IV-C Learning to land
In this task, the quadrotor starts in flight, gradually decelerates, and eventually lands safely on the landing area. It continuously monitors the limited landing area directly below, which is represented as a black rectangle positioned horizontally on the ground, as shown in Figure 6. The quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in its field of view via threshold segmentation, strives to keep it centered, and incorporates it into its observation. During the flight, the quadrotor captures only $64\times 64$ RGB image from the camera rather than acquiring precise spatial information about the landing area. This training process takes 1.7 hours, whose extra time (compared to Navigation) accounts for computing the center of the landing area.
V Conclusion And Discussion
@@
VisFly does not build in high-level controllers like Model Predictive Control (MPC) because of its requirements for substantial computational resources. Currently research [6] has demonstrated that these basic controllers are enough to train autonomous policies. In future work, we plan to integrate real-world dynamics into VisFly to enable Hardware-in-the-Loop (HITL) capabilities, and keep updating it.